OpenStudy (anonymous):

What is the error and correction in finding the difference? 3/4 - 9/2 = 3 - 9/4 - 2 = (-6/2) = -3

OpenStudy (anonymous):

Help?

OpenStudy (aaliyahfairgood):

first find the lcm/ least common denominator

OpenStudy (aaliyahfairgood):

multiple, sorry

OpenStudy (aaliyahfairgood):

I mean lcd. so multiply all numbers by 8 3/4 - 9/2 = 3 - 9/4 - 2 = (-6/2) = -3 6-36 =24 - 18- 16 = 24 = -24